Saddle Inc. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the company’s operations.

Standard

Custom

Direct labor costs $50,000 $100,000
Machine hours 1,200 1,500
Setup hours 120 360


Total estimated overhead costs are $294,600. Overhead cost allocated to the machining activity cost pool is $189,000, and $105,600 is allocated to the machine setup activity cost pool.

Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e.g. 12.25.)
Predetermined overhead rate % of direct labor cost

LINK TO TEXT

LINK TO TEXT

Compute the overhead rates using the activity-based costing approach.
Machining $ per machine hour
Machine setup $ per setup hour

LINK TO TEXT

LINK TO TEXT

Determine the difference in allocation between the two approaches.
Traditional costing
Standard $
Custom $
Activity-based costing
Standard $
Custom $
Answer #1

1) Predetermine overhead rate = 294600/150000*100 = 196.40% of labor cost

2) Machining = 189000/2700 = 70 per MH

Machine setup = 105600/480 = 220 per setup

3) Allocation

Traditional costing
Standard (50000*1.964) 98200
Custom 196400
Activity based costing
Standard (1200*70+120*220) 110400
Custom (1500*70+360*220) 184200